# Detect, Explain, Escalate: Sustainable Dialogue Breakdown Management for LLM Agents

### v3

@@ V-A Dialogue Breakdown Datasets
Dialogue Breakdown Detection Challenge 5 (DBDC5). DBDC5, introduced at the WOCHAT IWSDS 2020 workshop, is a standard benchmark for dialogue breakdown detection [Higashinaka2020DBDC5, higashinaka_dialogue_2016]. The evaluation set includes 1,950 English and 2,672 Japanese dialogues between humans and conversational agents. Each system utterance is annotated by 15–30 human annotators into three categories: Breakdown, Non-Breakdown, and Possible Breakdown. The agents span rule-based, retrieval-based, and neural generative architectures. Although these differ from modern LLM-based systems, the underlying breakdown phenomena (topic drift, contradiction, ignored context) remain salient sources of error [Huang2025Taxonomy].
Breakdown Expectation for Task-Oriented Long Dialogues (BETOLD). Introduced in 2022, BETOLD targets breakdowns in real-world, task-oriented customer-service calls [terragni2022_BETOLD]. It contains 13,524 human–agent phone dialogues, annotated according to whether the interaction ends with a “late user-initiated hang-up or forward” (LUHF), indicating user frustration and either call termination or escalation. Around 33% of dialogues are labelled as breakdowns. To protect privacy, BETOLD encodes dialogues as sequences of intents and entities rather than raw text. We follow the official train–test splits and treat the provided labels as ground truth.
@@ V-B Data Preprocessing and Experimental Setup
We binarize labels on DBDC5 and BETOLD, merging Possible Breakdown with Breakdown, following [Ghassel2024Are, lin_semi-supervised_2022]. To accommodate privacy constraints, the BETOLD dataset utilizes intent and entity abstractions rather than raw utterances, potentially limiting the efficacy of specific prompting techniques. We parse these structured dialogues and present them in a text form like “System: Intent: X — Entities: Y” for each turn, thereby preserving the sequence and dialogue flow in anonymized form. Basic cleaning (removal of extraneous symbols, ensuring consistent turn indexing) is applied for both datasets. No additional data augmentation is used beyond what is provided in the datasets.
@@
For LLM-based experiments we require models to output three fields: a justification, a decision, and a confidence score. We found that the order of these outputs affects behaviour: prompting for a justification first and a decision second encourages more deliberative reasoning and leads to better calibration than requesting the label upfront. For AR and CL+AR, we restrict evaluation to 10% subsamples of each dataset due to high token cost. Preliminary experiments indicate that a two-pass analogical pipeline (analogy generation followed by decision) performs comparably to single-pass prompting, aligning with existing literature [analogical].
@@ V-C Fine-tuning Parameters
We fine-tune Llama-3.1 8B using Low-Rank Adaptation (LoRA, rank = 16) [hu2022lora]. We use AdamW with 8-bit parameters [adamw], a learning rate of $2\times 10^{-4}$, linear decay, batch size 8, weight decay 0.01, and train for three epochs on a single NVIDIA A100 40GB GPU. We hypothesize that a carefully fine-tuned small LLM can provide competitive accuracy while remaining suitable for real-time monitoring. Our trained dialogue disruption monitor is released on HuggingFace [^1] https://huggingface.co/aghassel/dialogue_disruption_monitor. Because LoRA updates only a subset of parameters, this training regime substantially reduces compute and energy consumption compared to full-parameter fine-tuning.
@@ V-D Evaluation Metrics
We report accuracy and F1 scores for both Breakdown (B) and Non-Breakdown (NB) classes. Accuracy measures overall correctness; class-specific F1 scores capture performance on the minority Breakdown class without ignoring the majority NB class. Following DBDC5 challenge practice [ng_improving_2020, lin_semi-supervised_2022], we report accuracy and per-class F1 on held-out evaluation sets (DBDC5 official evaluation and reserved BETOLD test splits).
@@ V-E Inference and Cost Considerations
We evaluate a broad set of LLMs, summarized in Table II. Proprietary models include OpenAI’s GPT-3.5 Turbo and GPT-4o [openai_gpt-4_2023], and Anthropic’s Claude-3.5 Haiku and Claude-3.5 Sonnet [ClaudeModelCard]. Open-source models include Mistral’s Mixtral 8x7B and Mixtral 8x22B [jiang_mixtral_2024], DeepSeek’s DeepSeek-R1 [deepseek], and Meta’s Llama-3.1 8B, Llama-3.3 70B, and Llama-3.1 405B [llama]. Inference is run via the OpenRouter API [^2] https://openrouter.ai with a temperature of 0. To ensure comparability across models, we cap total tokens at 2,048 per query, which also reflects practical constraints for longer AR/CL+AR prompts.
@@ VI-A Main Findings
Dialogue Disruption Monitor. Despite its size, our fine-tuned Llama-3.1 8B monitor achieves 81.5% accuracy on DBDC5 English, exceeding several larger models and approaching Llama-3.3 70B. On DBDC5 Japanese, it reaches 67.9% accuracy with balanced F1 scores, reflecting limited Japanese coverage in pretraining. On BETOLD, the monitor achieves 67.2% accuracy, a 7-point absolute improvement over the zero-shot Llama-3.1 8B baseline (60.2%). This demonstrates cross-dataset transfer of breakdown detection behaviour. On DBDC5 English, the monitor’s calibration (MSE = 4.9) is competitive with much larger models. To improve Japanese performance further, future work could explore models such as Llama-3.1 Swallow 8B, which are pre-trained on large-scale Japanese corpora [swallow:COLM2024].
@@ VI-B Impact of Prompting Strategies
AR and CL+AR on BETOLD. While AR and CL+AR techniques improved performance on natural dialogue datasets (DBDC5 tracks), their effectiveness declined significantly on BETOLD. Results (highlighted in grey in Table II) consistently showed lower performance. This degradation stems from BETOLD’s representation of dialogues through structured intents and entities rather than natural language utterances. This mismatch undermines generalization, as models were predominantly trained on conventional dialogue data. Error analysis revealed frequent instruction-following failures. Models either generated no analogous examples (instead directly solving the original dialogue) or produced overly brief analogies with weak alignment to the target dialogue. Even GPT-4o struggled with reliable compliance (Figure 5). BETOLD’s longer dialogues (20–30 turns vs. 10–20 turns in DBDC5) exacerbate these issues. Three analogous examples plus the original conversation frequently exhausted the token budget (Figure 4). Smaller variants (Llama-3.1 8B, Mixtral 8x7B, Claude-3.5 Haiku) showed poor analogy generation, likely due to their reduced capacity and limited training.
@@ VI-D Sensitivity Analysis
Calibration Regime ($T>0.5$). Beyond $T=0.5$, the monitor begins escalating low-confidence non-breakdown predictions, yielding coupled increases in safety and cost. This reflects an effective uncertainty safety net: raising $T$ recovers borderline breakdowns that would otherwise be missed. On the out-of-domain BETOLD set, increasing $T$ from 0.5 to 0.8 boosts safety from roughly 70% to $90\%$, at the expense of higher escalation.
@@ VI-E End-to-End Repair Capabilities
To address the need for closed-loop remediation, we evaluated the effectiveness of the ”Escalate” module in repairing detected breakdowns. We sampled 100 correctly detected breakdowns from the DBDC5 English test set. We passed the dialogue history and the monitor’s generated justification ($\hat{j}_{i}$) to the Superior Model (Claude-3.5 Sonnet) with instructions to repair the response.
Automated evaluation using an LLM-as-a-judge (configured to assess response validity) confirmed that the pipeline successfully resolved the breakdown in 97% of cases [Gu2024A]. Qualitative analysis reveals that the superior model effectively utilizes the monitor’s diagnostic context to generate coherent replacements. For example, when the assistant hallucinated a contradiction (”I do not play video games. I like to play video games”), the repair module used the monitor’s flag to generate a consistent, context-aware correction (”I do enjoy playing video games in my free time…”). This confirms that the pipeline effectively converts detection signals into actionable repairs.
@@ VI-F Resource Efficiency and Sustainability
Resource efficiency is a crucial factor influencing the practical deployment of language models. Advanced models such as Claude-3.5 Sonnet (400B parameters), DeepSeek-R1 (671B parameters), and GPT-4 (estimated 1.7 trillion parameters[epochFrontierLanguage]) demand substantial monetary costs, experience network-induced latency, and have a higher carbon footprint per query due to their considerable size and reliance on data-centre computation. For example, based on cloud-provider pricing from Amazon Web Services (AWS), querying Llama-3.1 70B is approximately 3.3 times more expensive than its 8B-parameter counterpart, while Llama-3.1 405B incurs a cost approximately 10.3 times higher [amazonbedrockpricing]. Note, this also does not consider the latency, which in practice also increases substantially with model size, network routing complexity, and prompt length. Additionally, using longer prompts (such as AR or CL+AR) significantly escalates costs due to the increased token usage per query.
In contrast, our fine-tuned Llama-3.1 8B monitor runs efficiently on a single A100 GPU, with latency under half a second per dialogue turn, making it suitable for per-turn monitoring. In our cost analysis (Appendix A), we estimate that a hierarchical deployment combining a 70B assistant, 405B escalations on 10% of turns, and an 8B monitor achieves a 54% cost reduction relative to running the entire dialogue on Llama-3.1 405B. Given that large deployments such as ChatGPT may consume on the order of 1,058.5 GWh annually [bestbrokersAIsPower], selectively invoking large models only when necessary can meaningfully reduce cost and environmental impact.
@@ VI-G Relevance of Breakdown Datasets to Modern LLMs
Although DBDC5 was developed for chat-oriented dialogue systems rather than modern LLM deployments, its breakdown labels and error categories continue to capture core failure modes that remain prevalent in current deployments [Higashinaka2020DBDC5]. DBDC5 explicitly models phenomena such as contradiction, wrong or missing information, ignored questions, and topic-transition errors, which map closely onto “intrinsic hallucinations” (conflicts with source context) and broader faithfulness errors in recent hallucination taxonomies [Huang2025Taxonomy, Higashinaka2020DBDC5]. Consequently, detecting these breakdown types remains a useful proxy for assessing LLM reliability on fundamental conversational errors, even as underlying architectures and training regimes evolve.
@@ VII Conclusion
Robustness in conversational AI is advancing along several dimensions: larger and better-trained LLMs, more targeted benchmarks such as DBDC5 and BETOLD, and reasoning techniques such as chain-of-thought, analogical prompting, and curriculum learning. Our work combines these elements into a practical framework for dialogue breakdown management. Empirically, we show that modern LLMs, both proprietary and open-source, achieve state-of-the-art performance on dialogue breakdown detection benchmarks. However, a gap remains to human-level performance, especially in handling complex and ambiguous conversations. Advanced prompting (e.g., CoT, AR) can yield further gains for some high-capacity models, but improvements are not uniform across datasets. Short, well-chosen exemplars often strike a better balance between performance and token usage. We find that eliciting both justifications and numeric confidence can reduce overconfidence and improve calibration, particularly when the model is prompted to justify before deciding. For high-volume, high-stakes applications, coupling a fast, fine-tuned monitor with on-demand escalation to a frontier LLM provides a practical path to reliability and sustainability. While closed-source models currently lead in absolute accuracy, careful tuning of small open-source models narrows the gap. Future work should explore more fine-grained breakdown taxonomies, richer multimodal settings, and tighter integration of monitoring signals into policy-level decisions, with the goal of building robust, interpretable, and resource-aware conversational systems at scale.
@@ VIII Limitations
Although our approach performs competitively on both English and Japanese DBDC5 benchmarks, the smaller Llama-3.1 8B model shows limited Japanese coverage, suggesting that stronger multilingual pretraining or dedicated Japanese adaptation (e.g., using Swallow [swallow:COLM2024]) is needed. Moreover, real-world dialogues involving code-switching, adversarial inputs, or highly specialized domains may demand additional adaptations beyond the benchmarks we evaluate. Second, prompt engineering remains model- and dataset-specific, and methods such as AR and CL+AR can be brittle and expensive. Moreover, LLMs remain partially opaque: even with chain-of-thought or analogical prompting, internal reasoning processes are not fully observable. Escalating to larger models introduces latency and energy overhead, which may be problematic in ultra-low-latency or resource-constrained settings. Finally, our binary breakdown labels elide nuanced distinctions between mild and severe breakdowns and do not explicitly capture repair quality or user satisfaction. More expressive annotation schemes and joint modelling of detection and repair could enable finer-grained control in deployment. We leave these directions to future work.
@@ Appendix A Cost-Savings Example on AWS Bedrock
Table III lists on-demand prices [^3] https://aws.amazon.com/bedrock/pricing/ for Meta’s Llama-3.1 family. Bedrock bills input and output tokens separately, so the effective cost per 1k tokens is the sum of the two columns.
@@ A-B Baseline Costs
  • Always 405B  $9.6\times\mathdollar 0.00480=\mathdollar 0.0461$
  • Always 70B  $9.6\times\mathdollar 0.00144=\mathdollar 0.0138$
@@ A-D Adding a Lightweight Monitor
$9\,600\text{ tokens}\;\times\;\mathdollar 0.00044\;=\;\mathdollar 0.00422.$
@@
$\mathdollar 0.00422\;(\text{monitor})\;+\;\mathdollar 0.01705\;(\text{dialogue})\;=\;\mathdollar 0.02127.$
Selective escalation with monitoring cuts cost by $1-\tfrac{0.02127}{0.0461}\approx 54\%$ relative to running the entire conversation on Llama-3.1 405B, yet preserves the option to leverage the larger model when necessary.

### v4

@@ V-A Dialogue Breakdown Datasets
Dialogue Breakdown Detection Challenge 5 (DBDC5). DBDC5, introduced at the WOCHAT IWSDS 2020 workshop, is a standard benchmark for dialogue breakdown detection [24, 23]. The evaluation set includes 1,950 English and 2,672 Japanese dialogues between humans and conversational agents. Each system utterance is annotated by 15–30 human annotators into three categories: Breakdown, Non-Breakdown, and Possible Breakdown. The agents span rule-based, retrieval-based, and neural generative architectures. Although these differ from modern LLM-based systems, the underlying breakdown phenomena (topic drift, contradiction, ignored context) remain salient sources of error [27].
Breakdown Expectation for Task-Oriented Long Dialogues (BETOLD). Introduced in 2022, BETOLD targets breakdowns in real-world, task-oriented customer-service calls [43]. It contains 13,524 human–agent phone dialogues, annotated according to whether the interaction ends with a “late user-initiated hang-up or forward” (LUHF), indicating user frustration and either call termination or escalation. Around 33% of dialogues are labelled as breakdowns. To protect privacy, BETOLD encodes dialogues as sequences of intents and entities rather than raw text. We follow the official train–test splits and treat the provided labels as ground truth.
@@ V-B Data Preprocessing and Experimental Setup
We binarize labels on DBDC5 and BETOLD, merging Possible Breakdown with Breakdown, following [21, 31]. To accommodate privacy constraints, the BETOLD dataset utilizes intent and entity abstractions rather than raw utterances, potentially limiting the efficacy of specific prompting techniques. We parse these structured dialogues and present them in a text form like “System: Intent: X — Entities: Y” for each turn, thereby preserving the sequence and dialogue flow in anonymized form. Basic cleaning (removal of extraneous symbols, ensuring consistent turn indexing) is applied for both datasets. No additional data augmentation is used beyond what is provided in the datasets.
@@
For AR and CL+AR, we restrict evaluation to 10% subsamples of each dataset due to high token cost. Preliminary experiments indicate that a two-pass analogical pipeline (analogy generation followed by decision) performs comparably to single-pass prompting, aligning with existing literature [52].
@@ V-C Fine-tuning Parameters
We fine-tune Llama-3.1 8B using Low-Rank Adaptation (LoRA, rank = 16) [26]. We use AdamW with 8-bit parameters [33], a learning rate of $2\times 10^{-4}$, linear decay, batch size 8, weight decay 0.01, and train for three epochs on a single NVIDIA A100 40GB GPU. We hypothesize that a carefully fine-tuned small LLM can provide competitive accuracy while remaining suitable for real-time monitoring. Our trained dialogue disruption monitor is released on HuggingFace [^2] https://huggingface.co/aghassel/dialogue_disruption_monitor. Because LoRA updates only a subset of parameters, this training regime substantially reduces compute and energy consumption compared to full-parameter fine-tuning.
@@ V-D Evaluation Metrics
We report accuracy and F1 scores for both Breakdown (B) and Non-Breakdown (NB) classes. Accuracy measures overall correctness; class-specific F1 scores capture performance on the minority Breakdown class without ignoring the majority NB class. Following DBDC5 challenge practice [36, 31], we report accuracy and per-class F1 on held-out evaluation sets (DBDC5 official evaluation and reserved BETOLD test splits).
@@ V-E Inference and Cost Considerations
We evaluate a broad set of LLMs, summarized in Table II. Proprietary models include OpenAI’s GPT-3.5 Turbo and GPT-4o [38], and Anthropic’s Claude-3.5 Haiku and Claude-3.5 Sonnet [2]. Open-source models include Mistral’s Mixtral 8x7B and Mixtral 8x22B [28], DeepSeek’s DeepSeek-R1 [12], and Meta’s Llama-3.1 8B, Llama-3.3 70B, and Llama-3.1 405B [44]. Inference is run via the OpenRouter API [^3] https://openrouter.ai with a temperature of 0. To ensure comparability across models, we cap total tokens at 2,048 per query, which also reflects practical constraints for longer AR/CL+AR prompts.
@@ VI-A Main Findings
Dialogue Disruption Monitor. Despite its size, our fine-tuned Llama-3.1 8B monitor achieves 81.5% accuracy on DBDC5 English, exceeding several larger models and approaching Llama-3.3 70B. On DBDC5 Japanese, it reaches 67.9% accuracy with balanced F1 scores, reflecting limited Japanese coverage in pretraining. On BETOLD, the monitor achieves 67.2% accuracy, a 7-point absolute improvement over the zero-shot Llama-3.1 8B baseline (60.2%). This demonstrates cross-dataset transfer of breakdown detection behaviour. On DBDC5 English, the monitor’s calibration (MSE = 4.9) is competitive with much larger models. To improve Japanese performance further, future work could explore models such as Llama-3.1 Swallow 8B, which are pre-trained on large-scale Japanese corpora [19].
@@ VI-B Impact of Prompting Strategies
AR and CL+AR on BETOLD. While AR and CL+AR techniques improved performance on natural dialogue datasets (DBDC5 tracks), their effectiveness on BETOLD was mixed, with larger models showing degradation while some smaller models benefited (highlighted in grey in Table II). This degradation may stem from BETOLD’s representation of dialogues through structured intents and entities rather than natural language utterances. This mismatch undermines generalization, as models were predominantly trained on conventional dialogue data. Error analysis revealed frequent instruction-following failures. Models either generated no analogous examples (instead directly solving the original dialogue) or produced overly brief analogies with weak alignment to the target dialogue. Even GPT-4o struggled with reliable compliance (Figure 5). BETOLD’s longer dialogues (20–30 turns vs. 10–20 turns in DBDC5) exacerbate these issues. Three analogous examples plus the original conversation frequently exhausted the token budget (Figure 4). Smaller variants (Llama-3.1 8B, Mixtral 8x7B, Claude-3.5 Haiku) showed poor analogy generation, likely due to their reduced capacity and limited training.
@@ VI-D Sensitivity Analysis
Calibration Regime ($T>0.5$). Beyond $T=0.5$, the monitor begins escalating low-confidence non-breakdown predictions, yielding coupled increases in safety and cost. This reflects an effective uncertainty safety net: raising $T$ recovers borderline breakdowns that would otherwise be missed. On the BETOLD set, increasing $T$ from 0.5 to 0.8 boosts safety from roughly 70% to $90\%$, at the expense of higher escalation.
@@ VI-E End-to-End Repair Capabilities
To assess closed-loop remediation, we sampled 100 correctly detected breakdowns from the DBDC5 English test set. We passed the dialogue history and the monitor’s justification ($\hat{j}_{i}$) to the Superior Model (Claude-3.5 Sonnet) with the instruction: “The following response was flagged as a breakdown because: [Justification]. Please rewrite the response to be coherent and consistent with the dialogue history.”
An LLM-as-a-judge [22] evaluation confirmed that conditioning on these justifications resolved 97% of sampled breakdowns. In contrast, a baseline prompted with the generic instruction “Please rewrite the response to be coherent and consistent with the dialogue history” (without the monitor’s explanation) achieved a lower resolution rate of 92%. Qualitative analysis indicates that the superior model leverages specific details in the monitor’s flag, such as noting a direct contradiction with a previous turn, to generate targeted corrections rather than simply attempting a generic regeneration.
@@ VI-F Resource Efficiency and Sustainability
Resource efficiency is a crucial factor influencing the practical deployment of language models. Advanced models such as Claude-3.5 Sonnet (400B parameters), DeepSeek-R1 (671B parameters), and GPT-4 (estimated 1.7 trillion parameters[14]) demand substantial monetary costs, experience network-induced latency, and have a higher carbon footprint per query due to their considerable size and reliance on data-centre computation. For example, based on cloud-provider pricing from Amazon Web Services (AWS), querying Llama-3.1 70B is approximately 3.3 times more expensive than its 8B-parameter counterpart, while Llama-3.1 405B incurs a cost approximately 10.3 times higher [1]. Note, this also does not consider the latency, which in practice also increases substantially with model size, network routing complexity, and prompt length. Additionally, using longer prompts (such as AR or CL+AR) significantly escalates costs due to the increased token usage per query.
In contrast, our fine-tuned Llama-3.1 8B monitor runs efficiently on a single A100 GPU, with latency under half a second per dialogue turn, making it suitable for per-turn monitoring. In our cost analysis (Appendix A), we estimate that a hierarchical deployment combining a 70B assistant, 405B escalations on 10% of turns, and an 8B monitor achieves a 54% cost reduction relative to running the entire dialogue on Llama-3.1 405B. Given that large deployments such as ChatGPT may consume on the order of 1,058.5 GWh annually [25], invoking large models only when necessary can meaningfully reduce cost and environmental impact.
@@ VI-G Relevance of Breakdown Datasets to Modern LLMs
Although DBDC5 was developed for chat-oriented dialogue systems rather than modern LLM deployments, its breakdown labels and error categories continue to capture core failure modes that remain prevalent in current deployments [24]. DBDC5 explicitly models phenomena such as contradiction, wrong or missing information, ignored questions, and topic-transition errors, which map closely onto “intrinsic hallucinations” (conflicts with source context) and broader faithfulness errors in recent hallucination taxonomies [27, 24]. Consequently, detecting these breakdown types remains a useful proxy for assessing LLM reliability on fundamental conversational errors, even as underlying architectures and training regimes evolve.
@@ VII Conclusion
Robustness in conversational AI is advancing along several dimensions: larger and better-trained LLMs, more targeted benchmarks such as DBDC5 and BETOLD, and reasoning techniques such as chain-of-thought, analogical prompting, and curriculum learning. Our work combines these elements into a practical framework for dialogue breakdown management. Empirically, we show that modern LLMs, both proprietary and open-source, achieve state-of-the-art performance on dialogue breakdown detection benchmarks. However, a gap remains to human-level performance, especially in handling complex and ambiguous conversations. Advanced prompting (e.g., CoT, AR) can yield further gains for some high-capacity models, but improvements are not uniform across datasets. Short, well-chosen exemplars often strike a better balance between performance and token usage. We find that eliciting both justifications and numeric confidence can reduce overconfidence and improve calibration, particularly when the model is prompted to justify before deciding. For high-volume, high-stakes applications, coupling a fast, fine-tuned monitor with on-demand escalation to a frontier LLM provides a practical path to reliability and sustainability. While closed-source models currently lead in absolute accuracy, careful tuning of small open-source models narrows the gap. Future work should explore more fine-grained breakdown taxonomies, richer multimodal settings, and tighter integration of monitoring signals into policy-level decisions, with the goal of building interpretable, and resource-aware conversational systems at scale.
@@ VIII Limitations
Although our approach performs competitively on both English and Japanese DBDC5 benchmarks, the smaller Llama-3.1 8B model shows limited Japanese coverage, suggesting that stronger multilingual pretraining or dedicated Japanese adaptation (e.g., using Swallow [19]) is needed. Moreover, real-world dialogues involving code-switching, adversarial inputs, or highly specialized domains may demand additional adaptations beyond the benchmarks we evaluate. Second, prompt engineering remains model- and dataset-specific, and methods such as AR and CL+AR can be brittle and expensive. Moreover, LLMs remain partially opaque: even with chain-of-thought or analogical prompting, internal reasoning processes are not fully observable. Escalating to larger models introduces latency and energy overhead, which may be problematic in ultra-low-latency or resource-constrained settings. Finally, our binary breakdown labels elide nuanced distinctions between mild and severe breakdowns and do not explicitly capture repair quality or user satisfaction. More expressive annotation schemes and joint modelling of detection and repair could enable finer-grained control in deployment. We leave these directions to future work.
@@ Appendix A Cost-Savings Example on AWS Bedrock
Table III lists on-demand prices [^4] https://aws.amazon.com/bedrock/pricing/ for Meta’s Llama-3.1 family. Although Bedrock bills input and output tokens separately, the rates per 1k tokens are identical for these models; we therefore calculate costs using the total token count and the applicable rate.
@@ A-B Baseline Costs
  • Always 405B  $9.6\times\mathdollar 0.00240=\mathdollar 0.02304$
  • Always 70B  $9.6\times\mathdollar 0.00072=\mathdollar 0.00691$
@@ A-D Adding a Lightweight Monitor
$9\,600\text{ tokens}\;\times\;\mathdollar 0.00022\;\approx\;\mathdollar 0.00211.$
@@
$\mathdollar 0.00211\;(\text{monitor})\;+\;\mathdollar 0.00852\;(\text{dialogue})\;=\;\mathdollar 0.01063.$
Selective escalation with monitoring cuts cost by $1-\tfrac{0.01063}{0.02304}\approx 54\%$ relative to running the entire conversation on Llama-3.1 405B, yet preserves the option to leverage the larger model when necessary.